OpenStudy (anonymous):

The orbit of the moon about the earth is approximately circular with a mean radius of 3.84x10^8. It takes 27.3 days for the moon to complete one revolution about the Earth. Find the mean orbital speed of the moon

OpenStudy (shadowlegendx):

@oldrin.bataku Come help if you can

OpenStudy (anonymous):

given the presumption that the Earth's orbit is approximately circular with radius \(R=3.84\times10^8\) we know the circumference of said orbit is \(2\pi R\). if it takes \(\Delta T=27.3\text{ days}\) per revolution (i.e. one whole circumference) the orbital speed is then the ratio of the circumference to the elapsed time:$$\bar v=\frac{2\pi R}{\Delta T}=\dots$$